OpenStudy (anonymous):

1x + 30 = ? 5x + 18 = ?

OpenStudy (noseboy908):

are both equations equal to eachother? if not, then you have to have x to solve.

OpenStudy (campbell_st):

well it could be anything... I'm picking ? = 102

OpenStudy (anonymous):

1x + 30 = x = -30

OpenStudy (anonymous):

u have to find the value of x right???

OpenStudy (anonymous):

3

OpenStudy (anonymous):

X=3

OpenStudy (noseboy908):

put 3 in for x in each eqn. to solve, in that case.

OpenStudy (anonymous):

When x = 3; 1(3)+30 = ? 3+30 = ? Solve for "?" 5(3)+18 = ? 15 + 18 = ? Solve for " ? "